OpenStudy (secret-ninja):

A system of equations is shown below: 3x - y = 2 x + y = 6 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (anonymous):

3x - y = 2 x + y = 6 One y is negative and the other is positive, so they cancel each other out. We are left with: 4x = 8 Divide each side by 4 to isolate x. x = 2 <----- Answer. (: We can plug it in to find y as well. 3(2) - y = 2 6 - y = 2 -6 -6 -y = -4 y = 4 We can plug both in to check the answer. 3x - y = 2 x + y = 6 3(2) - 4 = 2 6 - 4 = 2 2 = 2 True. x + y = 6 2 + 4 = 6 6 = 6 True.

OpenStudy (secret-ninja):

Thank you! :D

OpenStudy (anonymous):

You're welcome! (: